OpenStudy (anonymous):

1. 2 logbq + 8 logbt (1 point) logb (q2 + t8) logb(q2t8) logb(qt2 + 8) (2 + 8) logb (q + t) 2. 4 log x – 6 log (x + 2) (1 point) 24 log log x4(x + 2)6 log x (x + 2)24 none of these

OpenStudy (callisto):

For the first question: \[2 \log_bq + 8 \log_bt \]You need to use the following log properties to simplify it: 1) \(a\log x = \log x^a\) 2) \(\log m + \log n = \log (mn)\) Example: \[2 \log_bq=log_b q^2\] So, try to do the first question. If you don't understand, ask.

OpenStudy (anonymous):

2 logbq + 8 logbt = logb(q2t8) ???

OpenStudy (callisto):

Yes!!

OpenStudy (anonymous):

Thank you! It was easier than I thought.

OpenStudy (anonymous):

I'm still confused about the second question.

OpenStudy (callisto):

Use the property \[a\log b = \log b^a\] and \[\log m - \log n = \log \frac{m}{n}\] to do it

OpenStudy (anonymous):

Thank you. I got it right =D

OpenStudy (callisto):

Welcome :)